At Niagara Falls, 6200 m3 per second of water falls down a height of 49 m.
(a) What is the rate (in watts) at which gravitational potential energy is dissipated by
the falling water?
1
Expert's answer
2018-11-06T15:37:09-0500

N=E_P/t=mgh/t=(ρV·gh)/t=(1000·6200·9.8·49)/1=2977.24·〖10〗^6 W.